# Overdetermined Steklov eigenvalue problems on compact surfaces

### v1

@@ Abstract.
We consider the overdetermined problem given by
$\Delta u=0\mbox{ in }\Omega,\quad\frac{\partial u}{\partial\nu}=\sigma_{1}u\mbox{ on }\partial\Omega,\quad|\nabla u|=\mbox{constant on }\partial\Omega,$
where $\Omega$ is a connected, orientable, compact Riemannian surface with smooth boundary $\partial\Omega$, and $\sigma_{1}$ is the first non-zero Steklov eigenvalue of $\Omega$. We give a complete characterization of $\Omega$ when the overdetermined problem is solvable.
Precisely, we prove that the overdetermined problem above admits a (nontrivial) solution if and only if $\Omega$ is conformally equivalent to either the flat unit disk or a flat cylinder $L,L]\times S^{1}$ for certain $L\geq L_{0}$ by a conformal transformation which (up to scaling) is an isometry on the boundary. In particular, we completely determine the compact domains in $\mathbb{R}^{2}$ or in $\mathbb{H}^{2}$ such that the overdetermined problem is solvable.
@@
Steklov eigenvalues and eigenfunctions; overdetermined problems; compact surfaces.
@@ 1. Introduction
$\begin{cases}\Delta u=-1&\mbox{ in }\Omega,\\ u=0&\mbox{ on }\partial\Omega,\\ \frac{\partial u}{\partial\nu}=\mbox{constant}&\mbox{ on }\partial\Omega\end{cases}$
@@
In this paper, we focus on overdetermined Steklov eigenvalue problems. Let $(M^{n},g)$ be an $n$-dimensional complete Riemannian manifold, $n\geq 2$. For a bounded domain $\Omega\subset M$ with smooth boundary $\partial\Omega$, the Steklov problem is to find real numbers $\sigma$ such that there exists a non-zero function $u$ on $\Omega$ satisfying
(1.1) $\Delta u=0$ in $\Omega$,
(1.2) $\frac{\partial u}{\partial\nu}=\sigma u$ on $\partial\Omega$,
where $\nu$ is the outward unit normal on $\partial\Omega$. We call such a $\sigma$ a Steklov eigenvalue, and $u$ the corresponding Steklov eigenfunction. It is well known that Steklov eigenvalues are discrete and satisfy
@@
(1.3) $|\nabla u|=c\mbox{ on }\partial\Omega$
@@
They proved
@@
We remark that the constant $c$ in (1.3) must be positive when $u$ is an eigenfunction corresponding to a non-zero Steklov eigenvalue. Otherwise, $u\equiv 0$ on $\partial\Omega$ by (1.2), and then $u\equiv 0$ in $\Omega$ by using (1.1), a contradiction.
A natural question is whether Theorem 1.1 can be generalized. It is known that the results fails for $\sigma_{k}(k\geq 2)$ or for $\Omega\subset\mathbb{R}^{n}(n\geq 3)$ (cf. [AM94, AM96]). Very recently, Lee and Seo [LS25] studied the surface case. Observing the condition (1.3) is equivalent to
(1.3’) $\operatorname{Hess}u(\nabla u)=\psi u\mbox{ for some $\psi\in C^{\infty}(\partial\Omega)$ on }\partial\Omega,$
@@
In this paper, we study above overdetermined Steklov problem for arbitrary compact Riemannian surfaces, dropping both the simply connected condition in Theorem 1.1 and the restrictions on $K_{\Omega}$ and $\psi$ in Theorem 1.2.
For convenience, throughout this paper, all the surfaces (or domains) are assumed to be orientable with smooth boundary. We introduce some definitions and notations.
@@
A surface (or a domain) $\Omega$ is called solvable for certain overdetermined problem if it admits a (non-trivial) solution of this overdetermined problem. Otherwise, we say that $\Omega$ is unsolvable for this overdetermined problem.
@@
  • $\mathbb{A}_{L}=[-L,L]\times S^{1}$: a cylinder with the flat metric $\mathrm{d}t^{2}+\mathrm{d}\theta^{2}$, where $L>0$ and $S^{1}$ is the unit circle.
  • $L_{0}(\approx 1.2)$: the unique positive solution of the equation $1/x=\tanh x$;
@@
  • $R_{0}=2\arctan\tanh(L_{0}/2)\approx 0.99\approx 56.7^{\circ}$.
  • $(\Omega_{1},g_{1})\sim(\Omega_{2},g_{2})$: there exists a conformal transformation from $\Omega_{1}$ to $\Omega_{2}$ which (after rescaling $g_{1}$) is an isometry on the boundary.
Now our main result is stated as follows.
@@
Let $(\Omega,g_{\Omega})$ be a connected compact surface with boundary $\partial\Omega$. Then $\Omega$ is solvable for (OS-I) if and only if either $(\Omega,g_{\Omega})\sim\mathbb{D}$ or $(\Omega,g_{\Omega})\sim\mathbb{A}_{L}$ for certain $L\geq L_{0}$.
@@
From Example 3.4 and Example 4.2, we can further know that, if $\Omega\sim\mathbb{D}$, then every first eigenfunction solves (OS-I); if $\Omega\sim\mathbb{A}_{L}$, then only the “linear eigenfunction” solves (OS-I). Hence, if we require that every function in the first Steklov eigenspace solves (OS-I), then the critical cylinder $\mathbb{A}_{L_{0}}$ must be excluded.
It is worth noticing that the critical cylinder $\mathbb{A}_{L_{0}}$ is equivalent (in the $\sim$ sense) to a so-called “critical” catenoid in [FS11], which is of geometric interest since it is a free boundary solution in the ball. We may call $\mathbb{A}_{L}$ a supercritical cylinder when $L\geq L_{0}$. The readers may refer [FS11]*Sect. 3 and 4 for details.
We will separate Theorem 1.4 into two parts: the simply connected case Theorem 3.1 and the multiply connected case Theorem 4.1. Furthermore, applying this theorem to domains of $\mathbb{M}_{K}^{2}$, we prove the following result (cf. Corollaries 3.5, 4.5 and 4.6).
@@
A connected domain $\Omega\subset\mathbb{S}^{2}$ with boundary $\partial\Omega$ is solvable for (OS-I) if and only if $\Omega$ is either a spherical cap, or $\Omega\sim A_{\mathrm{sph}}(R)$ for certain $R\in[R_{0},\pi/2)$.
@@
In Sect. 3, we solve the case of the simply connected surfaces, where Proposition 3.3 plays a crucial role. Although the auxiliary function used is the same as before, finer and novel calculations and analyses are done. This ultimately allows us to employ the Gauss–Bonnet formula and the Weinstock inequality and prove Theorem 3.1.
@@
In the last Sect. 5, we consider another overdetermined Steklov problem for simply connected domains in the hemisphere or in the hyperbolic space, and give a characterization of such domains (Theorem 5.2), where the Serrin-type theorem for domains of $\mathbb{S}^{2}$ and $\mathbb{H}^{2}$ is applied.
@@ 2.1. Normal exponential maps
(2.1) $g_{\Omega}=\mathrm{d}x^{2}+\lambda^{2}(x,y)\mathrm{d}y^{2},$
@@ 2.2. Basic properties of Steklov eigenfunctions
For a Steklov eigenfunction $u$ corresponding to $\sigma_{k}(k\geq 1)$, we always have
(2.6) $\sigma_{k}\int_{\partial\Omega}u=\int_{\partial\Omega}\frac{\partial u}{\partial\nu}=\int_{\Omega}\Delta u=0,$
(2.7) $\sigma_{k}\int_{\partial\Omega}u^{2}=\int_{\partial\Omega}u\frac{\partial u}{\partial\nu}=\int_{\Omega}|\nabla u|^{2}+\int_{\Omega}u\Delta u=\int_{\Omega}|\nabla u|^{2}.$
Suppose $u|_{\partial\Omega}=c$ is constant, then $c$ must be 0 from (2.6) since $\sigma_{k}>0$. Now (2.7) implies $|\nabla u|^{2}\equiv 0$ in $\Omega$. Hence, $u\equiv 0$ in $\Omega$, a contradiction. ∎
@@
$\Delta_{\tilde{g}}u=0\iff\Delta_{g}u=0\mbox{ in }\Omega,$
$\langle\nabla^{\tilde{g}}u,\nu_{\tilde{g}}\rangle_{\tilde{g}}=\rho_{0}^{-1}\langle\nabla^{g}u,\nu_{g}\rangle_{g}\mbox{ on }\partial\Omega,\quad|\nabla^{\tilde{g}}u|^{2}_{\tilde{g}}=\rho_{0}^{-2}|\nabla^{g}u|^{2}_{g}\mbox{ on }\partial\Omega,$
@@ 3. The case of simply connected surfaces
Following in [PP91, LS25], we introduce the auxiliary function $f=\log|\nabla u|^{2}$, where $u$ is any first Steklov eigenfunction of $\Omega$. Due to Lemma 2.2, $f$ is well-defined. We need to analyze $f$ carefully.
@@
For any first Steklov eigenfunction of $\Omega$, the Laplacian of $f=\log|\nabla u|^{2}$ satisfies
$\frac{1}{2}\Delta f=K\mbox{ in }\Omega,$
@@
Choose a local orthonormal frame $\{e_{1},e_{2}\}$ on $\Omega$, and denote $\nabla u=u_{1}e_{1}+u_{2}e_{2},\,\operatorname{Hess}u(e_{i},e_{j})=u_{ij}$. Without loss of generality, we can assume $(u_{ij})$ is diagonal. Hence, we have $u_{11}+u_{22}=\Delta u=0,u_{12}=u_{21}=0$. The Ricci identity gives
@@
$\frac{1}{2}\Delta f$ $=\frac{1}{2}\frac{\Delta|\nabla u|^{2}}{|\nabla u|^{2}}-\frac{1}{2}\frac{|\nabla|\nabla u|^{2}|^{2}}{|\nabla u|^{4}}$
@@
In the lemma above, $u$ need not satisfy (1.3). However, we need this additional condition in the key proposition below.
@@
Let $\Omega$ be a simply connected surface and $u$ be a first Steklov eigenfunction satisfying (1.3). Then the normal derivative of $f=\log|\nabla u|^{2}$ on the boundary satisfies
$\frac{1}{2}\frac{\partial f}{\partial\nu}=\sigma_{1}-\kappa\mbox{ on }\partial\Omega,$
@@
Denote $F=\{p\in\partial\Omega\mid\frac{1}{2}\frac{\partial f}{\partial\nu}=\sigma_{1}-\kappa\mbox{ at }p\}$. It is sufficient to prove $F=\partial\Omega$.
Clearly, $F$ is closed while its complement set $\FF^{c}=\partial\Omega\backslash F$ is open. Consider the open set $G=\{p\in\partial\Omega\mid u_{y}(0,y)\neq 0\mbox{ at }p=\gamma(y)\}$. We make the following assertions.
@@
Claim 3: if $F^{c}$ is not empty, then $\frac{1}{2}\frac{\partial f}{\partial\nu}=-\kappa$ on $F^{c}$.
By Claims 1 and 2, $F$ is not empty. If $F^{c}$ is not empty, then $F^{c}$ is not closed, otherwise $F^{c}=\partial\Omega$ by the connectedness of $\partial\Omega$, which contradicts $F\neq\emptyset$. Hence, there is a point $p\in\overline{F^{c}}\cap F$. By Claim 3 and the continuity, we have $\frac{1}{2}\frac{\partial f}{\partial\nu}=-\kappa$ on $F^{c}$ at $p$ ; on the other hand, $\frac{1}{2}\frac{\partial f}{\partial\nu}=\sigma_{1}-\kappa$ at $p$ by the definition of $F$. This is impossible. Therefore, $F^{c}$ must be empty, i.e., $F=\partial\Omega$.
@@
(3.1) $0=\Delta u=u_{xx}+\frac{1}{\lambda^{2}}u_{yy}+\frac{\lambda_{x}}{\lambda}u_{x}-\frac{\lambda_{y}}{\lambda^{3}}u_{y}.$
@@
(3.7) $\frac{1}{2}\frac{\partial f}{\partial\nu}=-\frac{1}{2}f_{x}|_{x=0}=-\left.\frac{1}{2}\frac{|\nabla u|^{2}_{x}}{|\nabla u|^{2}}\right|_{x=0}.$
@@
(3.12) $\frac{1}{2}\frac{\partial f}{\partial\nu}=\sigma_{1}-\kappa.$
@@
(3.13) $\frac{1}{2}|\nabla u|^{2}_{x}\Big|_{x=0}=-\lambda_{x}u_{x}^{2}\Big|_{x=0}=\kappa|\nabla u|^{2}.$
Inserting this into (3.7), we obtain $\frac{1}{2}\frac{\partial f}{\partial\nu}=-\kappa$ on $F^{c}$. Claim 3 is proved. ∎
@@
$\int_{\Omega}K$ $=\int_{\Omega}\frac{1}{2}\Delta f=\int_{\partial\Omega}\frac{1}{2}\frac{\partial f}{\partial\nu}$
@@
In fact, the solvability of $D_{K}(R)$ can also be directly verified via the explicit expression of the first Steklov eigenspace of $D_{K}(R)$ (cf. [BS14, Xio22])
Consider the equality case in the Weinstock inequality for domains of space forms, we immediately obtain the following

### v2

@@ Abstract.
We investigate the overdetermined problem given by
$\Delta u=0\text{ in }\Omega,\quad\frac{\partial u}{\partial\nu}=\sigma_{1}u\text{ on }\partial\Omega,\quad|\nabla u|=\text{constant on }\partial\Omega,$
where $\Omega$ is a connected compact Riemannian surface with smooth boundary $\partial\Omega$, and $\sigma_{1}$ is the first nonzero Steklov eigenvalue of $\Omega$. We prove that this overdetermined problem admits a nontrivial solution if and only if $\Omega$ is $\sigma$-homothetic to either the flat unit disk or a flat cylinder $T,T]\times S^{1}$ for some $T\geq T_{1}$. This gives a complete answer to the question raised by Payne and Philippin in [Z. Angew. Math. Phys. 42(6), 864–873, 1991] for $\sigma=\sigma_{1}$ and arbitrary surfaces. In particular, we completely characterize compact domains in 2-dimensional space forms for which the overdetermined problem is solvable.
@@
Steklov eigenvalues and eigenfunctions; overdetermined problems; compact surfaces; harmonic functions.
@@ 1. Introduction
$\left\{\begin{aligned} \Delta u&=-1&&\text{ in }\Omega,\\ u&=0&&\text{ on }\partial\Omega,\\ \partial_{\nu}u&=\text{constant}&&\text{ on }\partial\Omega\end{aligned}\right$.
@@
This paper is devoted to overdetermined Steklov eigenvalue problems. Let $(M^{n},g)$ be an $n$-dimensional complete Riemannian manifold, $n\geq 2$. For a bounded domain $\Omega\subset M$ with smooth boundary $\partial\Omega$, the Steklov problem is to find real numbers $\sigma$ such that there exists a nonzero function $u$ on $\Omega$ satisfying
(1.1) $\Delta u$ $=0$ $\text{ in }\Omega,$
(1.2) $\partial_{\nu}u$ $=\sigma u\quad$ $\text{ on }\partial\Omega,$
where $\nu$ is the outward unit normal to $\partial\Omega$. We call such a $\sigma$ a Steklov eigenvalue, and $u$ the corresponding Steklov eigenfunction. It is well known that Steklov eigenvalues are discrete and satisfy
@@
(1.3) $|\nabla u|=c\text{ on }\partial\Omega$
@@
In the same paper, they studied the above overdetermined problem for planar domains and proved the following theorem.
@@
We remark that the constant $c$ in (1.3) must be positive when $u$ is an eigenfunction corresponding to a nonzero Steklov eigenvalue. Otherwise, $u\equiv 0$ on $\partial\Omega$ by (1.2), and then $u\equiv 0$ in $\Omega$ by (1.1), a contradiction.
A natural question is whether Theorem 1.2 can be generalized. In [PP91], Payne and Philippin conjectured that Theorem 1.2 holds for $\sigma_{k}(k\geq 2)$ and $\Omega\subset\mathbb{R}^{n}(n\geq 3)$. However, this conjecture turned out to be false. For planar domains, Alessandrini and Magnanini [AM94] established a relation between a solution $u$ and the associated holomorphic function $F$ (called “complex potential”). By analyzing the zeros of $F$ in $\Omega$, they constructed various non-symmetric solvable domains (cf. Definition 1.4). For $\Omega\subset\mathbb{R}^{n}(n\geq 3)$, Alessandrini and Magnanini [AM96] connected solutions to isoparametric functions and isoparametric hypersurfaces, and proved an analogue of Theorem 1.2 under additional conditions.
Very recently, Lee and Seo [LS25] studied the surface case. Observing the condition (1.3) is equivalent to
$\operatorname{Hess}u(\nabla u)=\psi\nu\text{ for some $\psi\in C^{\infty}(\partial\Omega)$ on }\partial\Omega,$
@@
We study above overdetermined problem for arbitrary compact Riemannian surfaces, dropping both the simply connected condition in Theorem 1.2 and the restrictions on $K_{\Omega}$ and $\psi$ in Theorem 1.3. Due to the examples constructed by Alessandrini and Magnanini [AM94]*Theorem 1.2, a complete classification of solvable surfaces seems elusive; therefore, we focus on the case $\sigma=\sigma_{1}$, and provide a further discussion for general $\sigma$.
For convenience, throughout this paper, all the surfaces (or domains) and there boundaries are assumed to be smooth. We introduce some definitions and notations.
@@
A surface (or a domain) $\Omega$ is called solvable for an overdetermined problem if it admits a nontrivial solution of this overdetermined problem. Otherwise, we say that $\Omega$ is unsolvable for this overdetermined problem.
@@
Under conditions on the Steklov spectrum, as pointed out in [FS16], one can only hope to classify surfaces up to the equivalence relation called $\sigma$-homothetic” (cf. Lemma 2.3 for details). This terminology was introduced by Fraser and Schoen.
Definition 1.5 (cf. [FS16]*Definition 2.1).
For two surfaces $(\Omega_{1},g_{1})$ and $(\Omega_{2},g_{2})$, we say that $\Omega_{1}$ and $\Omega_{2}$ are $\sigma$-homothetic if there is a conformal diffeomorphism $\Phi\colon\Omega_{1}\to\Omega_{2}$ such that the pullback metric $\Phi^{\ast}g_{2}=\rho^{2}g_{1}$ with $\rho|_{\partial\Omega_{1}}=c$ for a constant $c$. Such $\Phi$ is called a $\sigma$-homothety (from $\Omega_{1}$ to $\Omega_{2}$).
@@
  • $\mathbb{A}_{T}=[-T,T]\times S^{1}$: a cylinder with the flat metric $\mathrm{d}t^{2}+\mathrm{d}\theta^{2}$, where $L>0$ and $S^{1}$ is the unit circle.
  • $T_{1}(\approx 1.2)$: the unique positive solution of the equation $1/x=\tanh x$.
@@
  • $R_{1}=2\arctan\tanh(T_{1}/2)\approx 0.99\approx 56.7^{\circ}$.
  • $(\Omega_{1},g_{1})\sim(\Omega_{2},g_{2})$: $\Omega_{1}$ and $\Omega_{2}$ are $\sigma$-homothetic.
We now state our main result, which provides a complete answer to Question 1.1 for $\sigma=\sigma_{1}$ and any compact surfaces with smooth metrics.
@@
Let $(\Omega,g_{\Omega})$ be a connected compact surface with boundary $\partial\Omega$. Then $\Omega$ is solvable for (OS-I) if and only if $\Omega$ is $\sigma$-homothetic to either $\mathbb{D}$ or $\mathbb{A}_{T}$ for some $T\geq T_{1}$.
Furthermore, the solution set $\mathcal{S}$ of (OS-I) is a linear space and can be characterized as follows:
(1) For $(\Omega,g_{\Omega})\sim\mathbb{D}$, let $\Phi\colon\Omega\to\mathbb{D}$ be a $\sigma$-homothety. Then $\mathcal{S}$ coincides with the first Steklov eigenspace of $\Omega$, spanned by $\Phi^{\ast}(r\cos\theta)$ and $\Phi^{\ast}(r\sin\theta)$.
(2) For $(\Omega,g_{\Omega})\sim\mathbb{A}_{T},T\geq T_{1}$, let $\Phi\colon\Omega\to\mathbb{A}_{T}$ be a $\sigma$-homothety. Then $\mathcal{S}$ is the subspace of the first Steklov eigenspace of $\Omega$ spanned by $\Phi^{\ast}t$.
@@
(1) In this paper, we always treat the constant $c$ in (OS-I) as non-fixed; that is, we only require that $|\nabla u|$ be some constant on the boundary. If $c$ is prescribed, the solution is determined up to scaling of any nontrivial $u\in\mathcal{S}$.
(2) The second part of Theorem 1.6 follows directly from Example 3.4 and Example 4.2. Furthermore, we see that, if $\mathcal{S}$ coincides with the first Steklov eigenspace of $\Omega$, then the critical cylinder $\mathbb{A}_{T_{1}}$ must be excluded.
(3) It is worth noting that the critical cylinder $\mathbb{A}_{T_{1}}$ is $\sigma$-homothetic to a so-called “critical catenoid” in [FS11], which is of geometric interest since it is a free boundary solution in the ball, and it is the unique maximizer (up to $\sigma$-homothety) of $\sigma_{1}(\Omega)L(\partial\Omega)$ among all smooth annuli. The readers may refer to [FS11, FS16] for details. We may call $\mathbb{A}_{T}$ a supercritical cylinder when $T\geq T_{1}$.
We will separate Theorem 1.6 into two parts: the simply connected case Theorem 3.1 and the multiply connected case Theorem 4.1. Furthermore, applying this theorem to domains of $\mathbb{M}_{K}^{2}$, we prove the following results (cf. Corollaries 3.5, 4.6 and 4.7).
@@
A connected domain $\Omega\subset\mathbb{S}^{2}$ with boundary $\partial\Omega$ is solvable for (OS-I) if and only if $\Omega$ is either a spherical cap or $\sigma$-homothetic to $A_{\mathrm{sph}}(R)$ for some $R\in[R_{1},\pi/2)$. In the latter case, $\Omega$ is symmetric with respect to a certain great circle of $\mathbb{S}^{2}$.
Corollary 1.10.
A connected domain $\Omega\subset\mathbb{S}^{2}_{+}$ with boundary $\partial\Omega$ is solvable for (OS-I) if and only if $\Omega$ is a spherical cap.
@@
In Sect. 3, we solve the case of the simply connected surfaces, where Proposition 3.3 plays a crucial role. Regrading an auxiliary function, finer and novel calculations and analyses are done. This ultimately allows us to employ the Gauss–Bonnet formula and the Weinstock inequality and prove Theorem 3.1.
@@
In Sect. 5, we provide further discussions on several extensions. First, by relaxing the condition that the boundary gradient magnitude be a global constant to being locally constant, we establish a classification for this weaker overdetermined problem. Second, for higher eigenvalues, under the additional assumption that the overdetermined problem admits a solution without critical points, we prove a classification result. Finally, we derive a necessary condition for the overdetermined problem to be solvable in the general case.
In the last Sect. 6, we investigate another type of overdetermined condition, namely, that the mean value of each nonconstant Steklov eigenfunction over the domain vanishes. We prove that, for any compact connected domain in the hemisphere or hyperbolic space of arbitrary dimension, this overdetermined condition holds if and only if the domain is a geodesic ball (Theorem 6.2), where Serrin-type theorems for $\mathbb{S}^{n}$ and $\mathbb{H}^{n}$ is applied.
@@ 2.1. Normal exponential maps
$g_{\Omega}=\mathrm{d}x^{2}+\lambda^{2}(x,y)\mathrm{d}y^{2},$
@@ 2.2. Basic properties of Steklov eigenfunctions
For any Steklov eigenfunction $u$ corresponding to $\sigma_{k}(k\geq 1)$, $u$ is orthogonal to the constant function $u_{0}=1$ in the $L^{2}(\partial\Omega)$ sense. If $u|_{\partial\Omega}=c$ is constant, then
$0=\int_{\partial\Omega}u=cL(\partial\Omega),$
which implies $c=0$. Hence, $u\equiv 0$ in $\Omega$ by the harmonicity of $u$, a contradiction. ∎
@@
$\Delta_{\tilde{g}}u=0\iff\Delta_{g}u=0\text{ in }\Omega,$
$\langle\nabla^{\tilde{g}}u,\nu_{\tilde{g}}\rangle_{\tilde{g}}=\rho_{0}^{-1}\langle\nabla^{g}u,\nu_{g}\rangle_{g}\text{ on }\partial\Omega,\quad|\nabla^{\tilde{g}}u|^{2}_{\tilde{g}}=\rho_{0}^{-2}|\nabla^{g}u|^{2}_{g}\text{ on }\partial\Omega,$
@@ 3. The case of simply connected surfaces
We introduce an auxiliary function $f=\log|\nabla u|$ (cf. [PP91, LS25]), where $u$ is any first Steklov eigenfunction of $\Omega$. Due to Lemma 2.2, $f$ is well-defined. We need to analyze $f$ carefully.
@@
Assume $u$ is a harmonic function with no critical points in $\Omega$. Then the Laplacian of $f=\log|\nabla u|$ satisfies
$\Delta f=K\text{ in }\Omega,$
@@
We choose a local orthonormal frame $\{e_{1},e_{2}\}$ on $\Omega$, and denote $\nabla u=u_{1}e_{1}+u_{2}e_{2}$, $\operatorname{Hess}u(e_{i},e_{j})=u_{ij}$. Without loss of generality, we can assume $(u_{ij})$ is diagonal. Hence, we have $u_{11}+u_{22}=\Delta u=0,u_{12}=u_{21}=0$. The Ricci identity gives
@@
$\Delta f$ $=\frac{1}{2}\log|\nabla u|^{2}=\frac{1}{2}\frac{\Delta|\nabla u|^{2}}{|\nabla u|^{2}}-\frac{1}{2}\frac{|\nabla|\nabla u|^{2}|^{2}}{|\nabla u|^{4}}$
@@
We don’t need (1.3) when applying the above lemma to the first Steklov eigenfunction $u$. However, this additional condition is necessary in the key proposition below.
@@
Let $\Omega$ be a simply connected surface and $u$ be a first Steklov eigenfunction satisfying (1.3). Then the normal derivative of $f=\log|\nabla u|$ on the boundary satisfies
$\frac{\partial f}{\partial\nu}=\sigma_{1}-\kappa\text{ on }\partial\Omega,$
@@
Denote $F=\{p\in\partial\Omega\mid\frac{\partial f}{\partial\nu}=\sigma_{1}-\kappa\text{ at }p\}$. It is sufficient to prove $F=\partial\Omega$.
Clearly, $F$ is closed while its complement set $\FF^{c}=\partial\Omega\backslash F$ is open. Consider the open set $G=\{p\in\partial\Omega\mid u_{y}(0,y)\neq 0\text{ at }p=\gamma(y)\}$. We make the following assertions.
@@
Claim 3: if $F^{c}$ is not empty, then $\frac{\partial f}{\partial\nu}=-\kappa$ on $F^{c}$.
By Claims 1 and 2, $F$ is not empty. If $F^{c}$ is not empty, then $F^{c}$ is not closed, otherwise $F^{c}=\partial\Omega$ by the connectedness of $\partial\Omega$, which contradicts $F\neq\emptyset$. Hence, there is a point $p\in\overline{F^{c}}\cap F$. By Claim 3 and the continuity, we have $\frac{\partial f}{\partial\nu}=-\kappa$ on $F^{c}$ at $p$ ; on the other hand, $\frac{\partial f}{\partial\nu}=\sigma_{1}-\kappa$ at $p$ by the definition of $F$. This is impossible. Therefore, $F^{c}$ must be empty, i.e., $F=\partial\Omega$.
@@
$0=\Delta u=u_{xx}+\frac{1}{\lambda^{2}}u_{yy}+\frac{\lambda_{x}}{\lambda}u_{x}-\frac{\lambda_{y}}{\lambda^{3}}u_{y}.$
@@
(3.6) $\frac{\partial f}{\partial\nu}=-f_{x}|_{x=0}=-\left.\frac{1}{2}\frac{|\nabla u|^{2}_{x}}{|\nabla u|^{2}}\right|_{x=0}.$
@@
$\frac{\partial f}{\partial\nu}=\sigma_{1}-\kappa.$
@@
$\frac{1}{2}|\nabla u|^{2}_{x}\Big|_{x=0}=-\lambda_{x}u_{x}^{2}\Big|_{x=0}=\kappa|\nabla u|^{2}.$
Inserting this into (3.6), we obtain $\frac{\partial f}{\partial\nu}=-\kappa$ on $F^{c}$. Claim 3 is proved. ∎
@@
$\int_{\Omega}K$ $=\int_{\Omega}\Delta f=\int_{\partial\Omega}\frac{\partial f}{\partial\nu}$
@@
In fact, the solvability of $D_{K}(R)$ can also be directly verified via the explicit expression of the first Steklov eigenspace of $D_{K}(R)$ (cf. [BS14, Xio22]).
By considering the equality case of the Weinstock inequality for domains in space forms, we immediately obtain the following